OpenStudy (anonymous):

Find the next three terms of the sequence. Then write a rule for the sequence. 648, 216, 72, 24

OpenStudy (anonymous):

looks like you are dividing by...

OpenStudy (slaaibak):

or multiplying by...

OpenStudy (saifoo.khan):

or one of those^

OpenStudy (anonymous):

lol

OpenStudy (lgbasallote):

they all seem divisible by....

OpenStudy (anonymous):

i said this a=648 r=/3 n=4 an=648/3^n-1 but it was wrong

OpenStudy (saifoo.khan):

r = 216/648 r = 1/3

OpenStudy (lgbasallote):

itshouldnt be 4 i mean n

OpenStudy (lgbasallote):

you are looking for the 5th term so n = 5

OpenStudy (saifoo.khan):

T = ar^(n-1)

OpenStudy (lgbasallote):

\[\large A_5 = 648 \times (\frac{1}{3})^{5 - 1}\]

OpenStudy (lgbasallote):

i use a....but it's just same with what the dude in red said

OpenStudy (anonymous):

saifoo = dude in red? ho ho ho

OpenStudy (saifoo.khan):

It's not same. there is a difference b/w A and T.

OpenStudy (saifoo.khan):

Haha @satellite73

OpenStudy (anonymous):

must have been the guy who came down my chimney last christmas that is why i got coal in my stocking

OpenStudy (lgbasallote):

saifoo khlaus

OpenStudy (saifoo.khan):

ROFL. Dont forget me.. i gave you gifts. :D

OpenStudy (saifoo.khan):

What does khlaus means? o_O

OpenStudy (lgbasallote):

i dont know...just to make it obvious that it was a pun

OpenStudy (saifoo.khan):

Lol.